# A Stochastic Genetic Interacting Particle Method for Reaction-Diffusion-Advection Equations

### v1

@@ 2.3.2 Reaction Integration
2:Updated density field $\{\hat{u}_{n+1}^{j}\}_{j=1}^{M^{d}}$, masses $\{m_{n+1}^{i}\}$
@@
6:Reaction Integration:
7:for each bin $j$ do
8:  if $r(u)$ has closed-form solution then
9:   $\hat{u}_{n+1}^{j}\leftarrow\text{ExactIntegration}(\hat{u}_{*}^{j},r,\Delta t)$
10:  else
11:   $\hat{u}_{n+1}^{j}\leftarrow\text{ImplicitIntegration}(\hat{u}_{*}^{j},r,\Delta t)$
12:Mass Update:
@@
2.4 Resampling Step
Following the reaction integration that produces the updated density field $\{\hat{u}_{n+1}^{j}\}$, we employ a resampling strategy to redistribute particles according to this new density distribution. This step ensures particle efficiency and adaptivity while maintaining mass conservation.
@@ 2.4 Resampling Step
$p_{j}=\frac{\hat{u}_{n+1}^{j}\cdot(\Delta x)^{d}}{M_{n+1}},\quad j=1,\dots,M^{d}$ (13)
where $M_{n+1}=\sum_{j=1}^{M^{d}}\hat{u}_{n+1}^{j}\cdot(\Delta x)^{d}$ is the total mass after reaction computed in Alg.2. These probabilities guide the redistribution of particles during the resampling step to match the updated density distribution.
New particle counts are generated from a multinomial distribution:
$(n_{1},\dots,n_{M^{d}})\sim\text{Multinomial}(N,p_{1},\dots,p_{M^{d}})$ (14)
The resampling process considers the current particle distribution to minimize statistical errors:
@@
1:Intermediate particle positions $\{\mathbf{X}_{*}^{i}\}_{i=1}^{N}$, updated density field $\{\hat{u}_{n+1}^{j}\}_{j=1}^{M^{d}}$
@@
4:Compute $p_{j}=\hat{u}_{n+1}^{j}\cdot(\Delta x)^{d}/M_{n+1}$ for $j=1,\dots,M^{d}$
5:Sample $(n_{1},\dots,n_{M^{d}})\sim\text{Multinomial}(N,p_{1},\dots,p_{M^{d}})$
@@
15:  else
16:   $\mathbf{X}_{\text{new},j}\leftarrow\text{UniformSample}(\text{bin}_{j},n_{j})$ $\triangleright$ Sample uniformly from empty bin
17:  Append $\mathbf{X}_{\text{new},j}$ to $\mathbf{X}_{\text{new}}$
@@
19:Set $m_{n+1}^{i}=M_{n+1}/N$ for all $i=1,\dots,N$
  • When $n_{j}\leq c_{j}$: Sample without replacement to preserve diversity
  • When $n_{j}>c_{j}$: Sample with replacement to meet the target count
  • When $c_{j}=0$: Generate new particles uniformly within the bin
This approach minimizes statistical errors and maintains good representation of the density field while adapting to the evolving solution structure.
@@
1:NN, $T$, $\Delta t$, $L$, $d$, $u_{0}(\mathbf{x})$, $\mathbf{v}(\mathbf{x},t)$, $D$, $r(u)$, $M$
@@
14:return $\{\hat{u}_{n}\}_{n=0}^{n_{T}}$
Each time step requires $\mathcal{O}(N)$ operations for particle motion, $\mathcal{O}(N)$ for bin assignment, and $\mathcal{O}(M^{d})$ for density estimation and resampling. The overall complexity per time step is $\mathcal{O}(N+M^{d})$, making the method efficient for moderate-dimensional problems.
Remark 1 (Fast adaptive algorithm for computational efficiency).
We note that for reaction-diffusion equations with FKPP type nonlinearities where the solution remains bounded ($u_{\text{max}}=1$), significant computational savings can be achieved. In future work, we will develop an adaptive acceleration technique that exploits the known propagation behavior of the equation. Specifically, regions where $u=1$ (saturated regions) can be identified and excluded from particle-based computation, particles are then only employed in the non-saturated transition regions where $u<1$. This adaptive approach dramatically reduces the number of active bins $M^{d}_{\text{active}}\ll M^{d}$ required for density estimation and resampling, leading to substantial speedup, particularly in higher-dimensional problems where the curse of dimensionality would otherwise limit practical computations.
@@ 3 Error Analysis of the SGIP method for RDA Equation
In this section, we present a rigorous error analysis for the SGIP algorithm applied to the RDA equation. The analysis systematically decomposes the total error into four fundamental components: splitting error, particle approximation error, spatial discretization error, and sampling error. We establish precise bounds for each component and combine them to obtain the overall error estimate.
Let $u(t)$ denote the exact solution of the FKPP equation, $u^{\mathrm{Lie}}(t)$ denote the exact Lie splitting solution (Eqs.(2)(3)), and $\widetilde{u}_{n}$ denote the numerical SGIP solution at time $t_{n}=n\Delta t$.
We analyze the local error of the SGIP by decomposing it into distinct components corresponding to different stages of the algorithm. The one-step error at time $t_{n+1}$ admits the following decomposition:
$\begin{aligned} \|\tilde{u}_{n+1}-u_{n+1}\|_{L^{2}}&\leq\underbrace{\|\tilde{u}_{n+1}-\tilde{u}_{n+1}^{\text{Lie}}\|_{L^{2}}}_{\varepsilon_{\text{num}}}+\underbrace{\|\tilde{u}_{n+1}^{\text{Lie}}-u_{n+1}^{\text{Lie}}\|_{L^{2}}}_{\varepsilon_{\text{propag}}}+\underbrace{\|u_{n+1}^{\text{Lie}}-u_{n+1}\|_{L^{2}}}_{\varepsilon_{\text{split}}}\end{aligned}$ (15)
where $u^{\mathrm{Lie}}_{n+1}=S_{\mathcal{R}}(\Delta t)S_{\mathcal{A}}(\Delta t)u_{n}$ is the Lie splitting solution from the previous exact solution $u_{n}$. Here, $\varepsilon_{\mathrm{split}}$ represents the splitting error due to operator splitting, while $\varepsilon_{\mathrm{num}}$ captures the numerical implementation error arising from the particle discretization, and $\varepsilon_{\mathrm{propag}}$ denotes the error from propagating the numerical solution over one time step using the splitting scheme.
The numerical implementation error $\varepsilon_{\mathrm{num}}$ can be further decomposed by tracing through the algorithmic stages:
$\varepsilon_{\mathrm{num}}$ $=\|\tilde{u}_{n+1}-\widetilde{u}^{\mathrm{Lie}}_{n+1}\|_{L^{2}}$
$\leq\underbrace{\|S_{\mathcal{R}}(\Delta t)(\tilde{u}^{\mathrm{D}}_{n+1})-S_{\mathcal{R}}(\Delta t)S_{\mathcal{A}}(\Delta t)\tilde{u}_{n}\|_{L^{2}}}_{\varepsilon_{\mathrm{dens}}}$
$\quad+\underbrace{\|\tilde{u}^{\mathrm{R}}_{n+1}-S_{\mathcal{R}}(\Delta t)(\tilde{u}^{\mathrm{D}}_{n+1})\|_{L^{2}}}_{\varepsilon_{\mathrm{react}}}$
$\quad+\underbrace{\|\tilde{u}_{n+1}-\tilde{u}^{\mathrm{R}}_{n+1}\|_{L^{2}}}_{\varepsilon_{\mathrm{resamp}}}$ (16)
where the intermediate quantities are:
  • $\widetilde{u}^{\mathrm{D}}_{n+1}$: numerical density after particle advection-diffusion and binning,
  • $\widetilde{u}^{\mathrm{R}}_{n+1}$: numerical density after reaction integration step,
  • $\widetilde{u}_{n+1}$: final numerical solution after resampling step.
The numerical implementation error comprises four distinct components: $\varepsilon_{\mathrm{dens}}$ represents the density estimation error from particle advection-diffusion and spatial binning; $\varepsilon_{\mathrm{react}}$ corresponds to the reaction integration error from numerically solving the reaction ODEs; and $\varepsilon_{\mathrm{resamp}}$ captures the resampling error introduced during particle redistribution.
To establish rigorous error bounds, we introduce the following technical assumptions:
@@
  2. 2.
    Velocity field regularity: The velocity field $\mathbf{v}(\mathbf{x},t)$ is uniformly Lipschitz continuous in space:
    $|\mathbf{v}(\mathbf{x},t)-\mathbf{v}(\mathbf{y},t)|\leq L_{v}|\mathbf{x}-\mathbf{y}|\quad\text{for all }\mathbf{x},\mathbf{y}\in\Omega,t\in[0,T],$ (17)
    and is uniformly bounded: $\|\mathbf{v}\|_{L^{\infty}(\Omega\times[0,T])}\leq M_{1}$.
  3. 3.
    Boundedness: The exact solution $u(t,\cdot)$ remains uniformly bounded in the interval $[0,1]$ for all $t\in[0,T]$, and maintains smoothness throughout the time evolution.
  4. 4.
    Semigroup contractivity: The boundary conditions (Neumann, periodic, or decay conditions on $\mathbb{R}^{d}$) are such that the advection-diffusion semigroup $S_{\mathcal{A}}(t)=e^{t\mathcal{A}}$ is contractive in $L^{2}(\Omega)$, i.e.,
    $\|S_{\mathcal{A}}(t)u\|_{L^{2}}\leq\|u\|_{L^{2}}\quad\text{for all }t\geq 0.$ (18)
  5. 5.
@@
Assumption 2 (CFL-like Condition).
In the SGIP algorithm, we assume that the bin size $\Delta x$ approaches 0 faster than the time step $\Delta t$ approaches 0. Additionally, the number of particles $N$ is assumed to grow sufficiently fast relative to $\Delta x$ and $\Delta t$.
More formally, as $\Delta t\to 0$, $\Delta x\to 0$, and $N\to\infty$, there exist parameters $\kappa,\nu$ such that:
$\kappa:=\frac{\Delta x}{\Delta t},\ \kappa\to 0,\qquad\nu:=\frac{1}{\sqrt{N\Delta x^{d}}\Delta t},\ \nu\to 0.$ (20)
@@
We first present the main convergence theorem, with its proof relying on the technical lemmas established in the subsequent subsection.
Theorem 3 (Global error bound).
@@
Under Assumptions 1 and 2, for $n\Delta t\leq T$, the global error of the SGIP algorithm satisfies:
$\sup_{0\leq n\leq T/\Delta t}\mathbb{E}\,\|\widetilde{u}_{n}-u_{n}\|_{L^{2}}\leq C_{T}\left(\Delta t+(N\Delta x^{d}\Delta t)^{-1/2}+\frac{\Delta x}{\Delta t}\right)$ (21)
[…]
@@
The proof follows from the subsequent lemmas that bound each error component.
Lemma 1 (Propagation error estimate).
@@
Under Assumption 1, the propagation error satisfies:
$\|S_{\mathcal{R}}(\Delta t)S_{\mathcal{A}}(\Delta t)\tilde{u}_{n}-S_{\mathcal{R}}(\Delta t)S_{\mathcal{A}}(\Delta t)u_{n}\|_{L^{2}}\leq(1+C_{1}\Delta t)\|\tilde{u}_{n}-u_{n}\|_{L^{2}}$ (22)
@@
Proof.
We analyze the propagation error by considering each operator separately.
For the advection-diffusion operator $\mathcal{A}u=-\nabla\cdot(\mathbf{v}u)+D\Delta u$, consider two solutions $u(t)=S_{\mathcal{A}}(t)u_{0}$ and $v(t)=S_{\mathcal{A}}(t)v_{0}$ with difference w=uvw=u-v. Then:
$\partial_{t}w$ $=\mathcal{A}w=-\nabla\cdot(\mathbf{v}w)+D\Delta w$ (23)
Taking the $L^{2}$ inner product with $w$:
$\frac{1}{2}\frac{d}{dt}\|w\|_{L^{2}}^{2}$ $=\langle w,-\nabla\cdot(\mathbf{v}w)\rangle+D\langle w,\Delta w\rangle$ (24)
For the advection term, using integration by parts and the boundedness of $\mathbf{v}$:

### v2

@@ 2.4 Resampling Step
2:Updated density field $\{\hat{u}_{n+1}^{j}\}_{j\in\mathcal{G}}$, masses $\{m_{n+1}^{i}\}$
@@
6:end for
7:Reaction Integration:
8:for each bin $j$ do
9:  if $r(u)$ has closed-form solution then
10:   $\hat{u}_{n+1}^{j}\leftarrow\text{ExactIntegration}(\hat{u}_{*}^{j},r,\Delta t)$
11:  else
12:   $\hat{u}_{n+1}^{j}\leftarrow\text{ImplicitIntegration}(\hat{u}_{*}^{j},r,\Delta t)$
13:  end if
14:end for
15:Mass Update:
@@
(12) $p_{j}=\frac{\hat{u}_{n+1}^{j}\cdot(\Delta x)^{d}}{M_{n+1}},\quad j\in\mathcal{G},$
where $M_{n+1}=\sum_{j\in\mathcal{G}}\hat{u}_{n+1}^{j}\cdot(\Delta x)^{d}$ is the total mass after reaction computed in Alg.2. This probabilistic framework introduces a selection pressure analogous to natural selection in genetics, in that particles are more likely to survive in a bin—and thus be propagated to the next generation—if that bin contains more mass (is “healthier”) than its neighbors. Beyond this biological analogy, resampling is critical for long-term numerical stability and computational efficiency, as it continuously rebalances the particle ensemble to control statistical variance.
To implement this redistribution, the target particle counts per bin are drawn from a multinomial distribution parameterized by these probabilities:
(13) $(n_{j})_{j\in\mathcal{G}}\sim\text{Multinomial}(N,(p_{j})_{j\in\mathcal{G}}).$
  • When $n_{j}\leq c_{j}$: Sample without replacement to preserve diversity
  • When $n_{j}>c_{j}$: Sample with replacement to meet the target count
  • When $c_{j}=0$: Generate new particles uniformly within the bin
Algorithm 3 implements the resampling strategy described above, adaptively redistributing particles according to the updated density field while preserving statistical accuracy. This approach minimizes statistical errors and maintains a good representation of the density field while adapting to the evolving solution structure.
@@
1:Intermediate particle positions $\{\mathbf{X}_{*}^{i}\}_{i=1}^{N}$, updated density field $\{\hat{u}_{n+1}^{j}\}_{j\in\mathcal{G}}$
@@
4:Compute $p_{j}=\hat{u}_{n+1}^{j}\cdot(\Delta x)^{d}/M_{n+1}$ for $j\in\mathcal{G}$
5:Sample $(n_{j})_{j\in\mathcal{G}}\sim\text{Multinomial}(N,(p_{j})_{j\in\mathcal{G}})$
@@
15:   end if
16:  else
17:   $\mathbf{X}_{\text{new},j}\leftarrow\text{UniformSample}(\text{bin}_{j},n_{j})$ $\triangleright$ Sample uniformly from empty bin
18:  end if
19:  Append $\mathbf{X}_{\text{new},j}$ to $\mathbf{X}_{\text{new}}$
20:end for
@@
22:Set $m_{n+1}^{i}=M_{n+1}/N$ for all $i=1,\dots,N$
@@
1:NN, $T$, $\Delta t$, $L$, $d$, $u_{0}(\mathbf{x})$, $\mathbf{v}(\mathbf{x},t)$, $D$, $r(u)$, $K$
@@
14:end for
15:return $\{\hat{u}_{n}\}_{n=0}^{n_{T}}$
Each time step requires $\mathcal{O}(N)$ operations for particle motion, $\mathcal{O}(N)$ for bin assignment, and $\mathcal{O}(|\mathcal{G}|)$ for density estimation and resampling. The overall complexity per time step is $\mathcal{O}(N+|\mathcal{G}|)$, making the method efficient for moderate-dimensional problems.
Remark 2.1 (Potential for adaptive acceleration).
For reaction-diffusion equations with FKPP-type nonlinearities where the solution remains bounded ($\max_{\mathbf{x}}u(\mathbf{x},t)=1$), significant computational savings can be achieved. In future work, we will develop an adaptive acceleration technique that exploits the known propagation behavior of the equation. Specifically, regions where $u=1$ (saturated regions) can be identified and excluded from particle-based computation, and particles are then only employed in the non-saturated transition regions where $u<1$. This adaptive approach dramatically reduces the number of active bins $|\mathcal{G}_{\text{active}}|\ll|\mathcal{G}|$ required for density estimation and resampling, leading to substantial speedup, particularly in higher-dimensional problems where the curse of dimensionality would otherwise limit practical computations.
@@ 3 Error Analysis of the SGIP method for RDA Equation
In this section, we present a complete error analysis for the SGIP algorithm applied to RDA equations. The analysis systematically decomposes the total error into three main categories: the consistency error arising from operator splitting and spatial discretization, the propagation error governed by the stability of the evolution operators, and the stochastic numerical error introduced by density estimation and resampling. We establish precise bounds for each component and combine them to obtain the overall error estimate.
We recall the operator splitting framework for the RDA equation introduced in Eqs.(2)-(3):
(14) $\frac{\partial u}{\partial t}=\mathcal{A}u+\mathcal{R}u,\quad\text{where}\quad\mathcal{A}u=D\Delta u-\nabla\cdot(\mathbf{v}u),\quad\mathcal{R}u=r(u).$
Let $\mathcal{S}_{\mathcal{A}}(t)$ and $\mathcal{S}_{\mathcal{R}}(t)$ denote the exact solution semigroups generated by $\mathcal{A}$ and $\mathcal{R}$, respectively. The ideal Lie-Trotter splitting evolution over a time step $\Delta t$ is given by $\mathcal{S}_{\mathcal{R}}(\Delta t)\circ\mathcal{S}_{\mathcal{A}}(\Delta t)$.
Let $u(t)$ denote the exact solution of the RDA equation, and $\widetilde{u}_{n}$ denote the numerical SGIP solution at time $t_{n}=n\Delta t$. To establish precise error bounds, we introduce the following technical assumption:
@@
  2. (b)
    Velocity field regularity: The velocity field $\mathbf{v}(\mathbf{x},t)$ is uniformly bounded and Lipschitz continuous in space. There exist constants $M_{1}$ and $L_{v}$ such that:
    (15) $\|\mathbf{v}\|_{L^{\infty}(\Omega\times[0,T])}\leq M_{1},\quad|\mathbf{v}(\mathbf{x},t)-\mathbf{v}(\mathbf{y},t)|\leq L_{v}|\mathbf{x}-\mathbf{y}|.$
  3. (c)
    Boundedness: There exists a constant $C_{u}>0$ such that the exact solution $u(t,\cdot)$ remains uniformly bounded in the interval $0,C_{u}]$ for all $t\in[0,T]$, and maintains smoothness throughout the time evolution. (For the FKPP equation, $C_{u}=1$ is the natural carrying capacity.)
  4. (d)
@@
  5. (e)
    Structural bound preservation: The reaction function $r(u)$ satisfies
    (17) $r(0)\geq 0\quad\text{and}\quad r(u)\leq 0\quad\text{for all }u\geq C_{u}.$
@@
We first present the main convergence theorem, with its proof relying on the technical lemmas established later in this section.
Theorem 3.1 (Global error bound).
@@
Under Assumption 1, for $n\Delta t\leq T$, the global error of the SGIP algorithm satisfies:
(18) $\sup_{0\leq n\leq T/\Delta t}\mathbb{E}\,\|\widetilde{u}_{n}-u_{n}\|_{L^{2}}\leq C_{T}\left(\Delta t+\frac{1}{\sqrt{N(\Delta x)^{d}\Delta t}}+\Delta x\right),$
where the constant $C_{T}$ depends on the final time $T$, the diffusion parameter $D$, the domain size $L$, the dimension $d$ and the solution regularity.
A direct consequence of Theorem 3.1 is that, under Assumption 1, the SGIP solution $\widetilde{u}_{n}$ converges to the exact solution $u_{n}$ in $L^{2}$-norm as $\Delta t,\Delta x\to 0$ and $N\to\infty$, provided that the scaling condition $N(\Delta x)^{d}\Delta t\to\infty$ is satisfied to ensure the vanishing of statistical variance.
Remark 3.2 (Practical convergence and scaling).
In practical simulations where $\Delta t$ and $\Delta x$ are typically fixed, the error bound (18) indicates that the statistical noise is the only term dependent on $N$. Crucially, as demonstrated by the numerical experiments in Section 4, high accuracy is already attained with moderate values of $\Delta x$, $\Delta t$, and $N$, which is far less restrictive than those suggested by the theoretical analysis. This mild dependence on discretization parameters and sample size enables both robustness and efficiency, even in challenging 3D settings.
Before presenting the detailed error estimates, we establish the boundedness of the expected total mass of the SGIP solution.
Lemma 3.3 (Boundedness of Total Mass).
Under Assumption 1, the expected total mass of the SGIP solution remains bounded over the finite time interval $[0,T]$. Specifically, there exists a constant $C_{M}>0$ such that for all $n\Delta t\leq T$:
(19) $\mathbb{E}[M_{n}]:=\mathbb{E}\left[\int_{\Omega}\widetilde{u}_{n}(\mathbf{x})d\mathbf{x}\right]\leq C_{M}.$
Proof 3.4.
The proof relies on the mass evolution properties of the algorithm. Let $M_{n}$ be the total mass at step $n$. Since the advection-diffusion step conserves mass exactly and the resampling step preserves mass in expectation, the evolution of the expected total mass is determined solely by the reaction dynamics. From Assumption 1(d)(e), we have $r(u)\leq C_{u}+L_{r}u$ for $u\geq 0$. Thus, the growth of total mass is bounded by a linear function of the mass itself:
(20) $\frac{d}{dt}M(t)=\int_{\Omega}r(u)d\mathbf{x}\leq\int_{\Omega}(C_{u}+L_{r}u)d\mathbf{x}=C_{u}|\Omega|+L_{r}M(t).$
By Gronwall’s inequality, the mass grows at most exponentially. Since we consider a finite time interval $[0,T]$, the expected mass $\mathbb{E}[M_{n}]$ remains uniformly bounded by a constant $C_{M}$ dependent on $T,\Omega,L_{r}$ and the initial mass $M_{0}$.
We now analyze the local error of the SGIP by decomposing it into distinct components corresponding to different stages of the algorithm. To distinguish between the accumulation of numerical errors and the irreducible spatial resolution error, we introduce the projection of the exact solution onto the piecewise constant finite volume grid. Let $\bar{u}_{n}=\mathcal{P}_{h}u(t_{n})$ denote the cell averages of the exact solution at time $t_{n}$. Unless otherwise stated, all subsequent norms $\|\cdot\|$ will refer to the $L^{2}$-norms.
We decompose the total error into dynamic and static parts:
(21) $\|\tilde{u}_{n}-u_{n}\|\leq\underbrace{\|\tilde{u}_{n}-\bar{u}_{n}\|}_{\varepsilon_{\text{dyn}}}+\underbrace{\|\bar{u}_{n}-u_{n}\|}_{\varepsilon_{\text{proj}}}.$
Here, $\varepsilon_{\text{proj}}$ represents the projection error, which is bounded by $O(\Delta x)$ for smooth solutions and does not accumulate over time. We focus our analysis on the evolution of the dynamic error $\varepsilon_{\text{dyn}}$. To define the numerical error precisely, we introduce the ideal discrete evolution operator $\Phi_{h}(u)=S_{\mathcal{R}}(\Delta t)\mathcal{P}_{h}S_{\mathcal{A}}(\Delta t)u$, which represents the expected one-step evolution of the algorithm excluding stochastic effects. At time $t_{n+1}$, the error can be decomposed as:
(22) $\begin{aligned} \|\tilde{u}_{n+1}\!-\!\bar{u}_{n+1}\|\!&\leq\!\underbrace{\|\tilde{u}_{n+1}\!-\!\Phi_{h}(\tilde{u}_{n})\|}_{\varepsilon_{\text{num}}}\!+\!\underbrace{\|\Phi_{h}(\tilde{u}_{n})\!-\!\Phi_{h}(\bar{u}_{n})\|}_{\varepsilon_{\text{propag}}}\!+\!\underbrace{\|\Phi_{h}(\bar{u}_{n})\!-\!\bar{u}_{n+1}\|}_{\varepsilon_{\text{consist}}}.\end{aligned}$
In this decomposition, $\varepsilon_{\mathrm{consist}}$ denotes the local consistency error, which measures the discrepancy between the ideal discrete scheme and the projection of the exact evolution. $\varepsilon_{\mathrm{propag}}$ represents the propagation error, reflecting the stability of the discrete operator acting on the error from the previous step. The term $\varepsilon_{\mathrm{num}}$ captures the stochastic numerical error, arising solely from the finite particle fluctuations around the ideal discrete operator.
The numerical implementation error $\varepsilon_{\mathrm{num}}$ can be further decomposed by tracing through the algorithmic stages. Let $\tilde{u}^{\mathrm{D}}_{n+1}$ denote the numerical density after particle advection-diffusion and binning, $\tilde{u}^{\mathrm{R}}_{n+1}$ the density after reaction integration, and $\tilde{u}_{n+1}$ the final numerical solution after resampling. Consistent with the definition of $\Phi_{h}$, the reference state for the density estimation is the projected advection $\mathcal{P}_{h}S_{\mathcal{A}}(\Delta t)\tilde{u}_{n}$. Then:
(23) $\varepsilon_{\mathrm{num}}$ $\leq\underbrace{\|S_{\mathcal{R}}(\Delta t)(\tilde{u}^{\mathrm{D}}_{n+1}-\mathcal{P}_{h}S_{\mathcal{A}}(\Delta t)\tilde{u}_{n})\|}_{\varepsilon_{\mathrm{dens}}}\!+\!\underbrace{\|\tilde{u}^{\mathrm{R}}_{n+1}-S_{\mathcal{R}}(\Delta t)(\tilde{u}^{\mathrm{D}}_{n+1})\|}_{\varepsilon_{\mathrm{react}}}\!+\!\underbrace{\|\tilde{u}_{n+1}-\tilde{u}^{\mathrm{R}}_{n+1}\|}_{\varepsilon_{\mathrm{resamp}}}.$
The three terms correspond to distinct error sources: $\varepsilon_{\mathrm{dens}}$ represents the statistical error of density estimation; $\varepsilon_{\mathrm{react}}$ corresponds to the ODE integration error; and $\varepsilon_{\mathrm{resamp}}$ captures the resampling noise introduced during particle redistribution.
@@
The proof of Theorem 3.1 follows from the subsequent lemmas that bound each error component in Eqs.(21)(22) and (23).
Lemma 3.5 (Propagation error estimate).
@@
Under Assumption 1, the propagation error of the ideal discrete operator $\Phi_{h}$ satisfies:
(24) $\|\Phi_{h}(\tilde{u}_{n})-\Phi_{h}(\bar{u}_{n})\|\leq(1+C_{1}\Delta t)\|\tilde{u}_{n}-\bar{u}_{n}\|,$
@@
Proof 3.6.
Let $w=\tilde{u}_{n}-\bar{u}_{n}$. We trace the evolution of the norm difference through the operators.
First, apply advection-diffusion: $w_{1}=S_{\mathcal{A}}(\Delta t)\tilde{u}_{n}-S_{\mathcal{A}}(\Delta t)\bar{u}_{n}$. Standard energy estimates for the advection-diffusion equation $\partial_{t}w_{1}=-\nabla\cdot(\mathbf{v}w_{1})+D\Delta w_{1}$ yield: